OpenStudy (anonymous):

Simply the complex rational expression: Show Work: (x/1 - x/x-5) / x-6 x x -- - --- 1 x-5 -------------- x-6

OpenStudy (zehanz):

Because \[\frac{ a }{ b }+\frac{ c }{ d }=\frac{ ad+bc }{ bd }\]and also:\[\frac{ \frac{ a }{ b } }{ c }=\frac{ a }{ b }\cdot\frac{ 1 }{ c }\]you can write:\[\frac{ x(x-5)-x }{ x-5 }\cdot \frac{ 1 }{ x-6 }\]

OpenStudy (anonymous):

wait so is the bottom part of the question the answer? sorry this question is my borthers hw so idk what to do!

OpenStudy (zehanz):

I partly simplified the expression. Now we have the product of two fractions: multiply the numerators and multiply the denominators to get one fraction:\[\frac{ x^2-5x-x }{ (x-5)(x-6) }=\frac{ x^2-6x }{ (x-5)(x-6) }\]This can be simplified further if you first factorize the numerator...

OpenStudy (anonymous):

can you just tell me the answer please :( this is the last question i will ask you hopefullly!!!!

OpenStudy (zehanz):

\[\frac{ x(x-6) }{ (x-5)(x-6) }=\frac{ x }{ x-5 }\]

OpenStudy (anonymous):

the choices for the answers in my packet are a) x-5 b) x/ x-5 c) x/ x-6 d) x-5/ x-6 @ZeHanz so would it be B

OpenStudy (zehanz):

yes, B, although I would write it as x/(x-5)

OpenStudy (anonymous):

ok thank you

OpenStudy (zehanz):

yw!